OpenStudy (anonymous):

if michael jordan has a vertical leap of 1.29 m, what is his take-off speed and his hang time (total time to move upwards to the peak and then return to the ground)?

OpenStudy (anonymous):

\[v ^{2}=u ^{2}+2as\] \[0=u ^{2}+ 2\times (-9.8) \times 1.29\] \[u=5.0283m/s \] ______________________________________________ \[v = u + at\] \[0 = 5.0283 -9.8 \times t\]\ \[t = 0.5130 \sec\] Total time = 2t = 1.0261 sec